OpenStudy (anonymous):

I am looking for the answer of : Write the standard form of the equation of the line passing through the point (-3, 5) and perpendicular to the line -4x -3y = -6

OpenStudy (amistre64):

looking for it? where did you last leave it?

OpenStudy (anonymous):

I pretty much got an answer but i need to know if it is right.

OpenStudy (amistre64):

type it up and all compare

OpenStudy (anonymous):

Does it have a slope of 3/4?

OpenStudy (amistre64):

I got mine written :)

OpenStudy (anonymous):

I have mine hidden under my hand so no one peeks :P

OpenStudy (anonymous):

:(... The answer i got was -4x + 3y = 27.

OpenStudy (anonymous):

hmmmm.... I ended up with -3x+4y=29

OpenStudy (amistre64):

-3x+4y = C -3 5 ----------- 9 + 20 = 29 -3x +4y = 29; or better yet 3x -4y = -29

OpenStudy (amistre64):

given ax + by = c slope = -a/b; so perp slope is b/a in other words; swap coeffs and flip a sign

OpenStudy (amistre64):

then calibrate the constant with the given point